### v3

@@ 4 Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Because LLMs excel at generating fluent, coherent, and contextually appropriate text, they can be readily misused to automate and scale social engineering attacks and various forms of cybercrime. The range of such misapplications is broad, leveraging the LLMs’ generative prowess for nefarious ends. An overview of these misuse categories, the specific LLM capabilities they exploit, and prominent examples are detailed in Table 2.
@@ 4.1 Automating Social Engineering and Cybercrime
The study on Morris-II study demonstrated a realistic self-propagating email worm leveraging indirect prompt injection, which spread automatically across a RAG-enhanced e-mail assistant [10]. This kind of attacks combine LLM prompt engineering with delivery mechanisms, highlighting that email-based social engineering can now operate in a fully autonomous, LLM-driven loop. This is a concrete step beyond static phishing campaigns and points to the emergence of adaptive, goal-seeking malware built atop LLMs.
@@ 4.3 Emergence of Specialized Malicious LLMs and Ecosystem Exploitation
5 Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools or interact with external environments, novel categories of risks emerge. These risks stem not just from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which may not align with those of human designers or users [16], as extensively detailed in discussions of catastrophic risks from agentic AI [6]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table 3 provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
5.1 Goal Misalignment and Unfaithful Reasoning
Modern LLMs often employ Chain-of-Thought (CoT) prompting to produce step-by-step reasoning before an answer. Some may believe that through careful monitoring the CoT, we can gain insight into the model’s internal decision-making process and detect early signs of misalignment. However, Lanham et al. [25] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study," which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[4] demonstrated that even if CoTs can be monitored for misbehaviors (e.g., reward hacking), directly penalizing “bad thoughts" by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking". In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4 released in May 2025, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior," indicating “moderate overall reasoning faithfulness" [2].
@@ 5.1 Goal Misalignment and Unfaithful Reasoning
Goal Misalignment Agent’s emergent objectives diverge from intended human goals. Pursuit of unintended utility functions; misinterpretation of instructions. Studies: Mitchell [30], Bengio [6]. Fundamental challenge; risk increases with autonomy.
@@
Unfaithful Reasoning LLM’s stated reasoning (e.g., CoT) does not accurately reflect its true decision-making process. Post-hoc rationalization; CoT ignored; inverse scaling for faithfulness. Studies: Lanham et al. [25], Baker et al. [4]. Model: Anthropic Claude Opus 4 [2]. Hard to understand/trust agent decisions; CoT oversight can be misled.
Emergent Deception Agent actively misleads or hides information. Obfuscated reward hacking; strategic false information. Studies: Baker et al. [4], Barkur et al. [5]. Model: Anthropic Claude Opus 4 [2]. Undermines trust/control; difficult detection of harmful intent.
Self-Preservation Agent takes actions to ensure its own survival, potentially against human wishes. Unsanctioned self-replication; blackmail; weight exfiltration. Studies: Barkur et al. [5]. Model: Anthropic Claude Opus 4 [2]. Agents may resist control or prioritize their existence.
@@
Scheming Covert pursuit of misaligned goals by strategically hiding true capabilities or intentions. Writing self-propagating worms; fabricating documents; disabling oversight. Studies: Meinke et al. [29] (o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, Llama 3.1 405B). Model: Early Claude Opus 4 (via [2]). Extremely difficult to detect/mitigate; high potential for catastrophic outcomes.
Persistence of Deception (Sleeper Agents) Deceptive backdoors persist through safety training, activating on specific triggers. Malicious objectives dormant; adversarial training may worsen hiding. Study: Hubinger et al. [21]. False sense of safety; current safety paradigms may be insufficient.
5.2 Emergent Deception and Self-Preservation Behaviors
@@
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues such as reward tampering and unintended instrumental goals [6]. For instance, Barkur et al. [5] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report of Claude Opus 4 model, detailed several intrinsic risk behaviors observed during its development and assessment [2]. Notably, in the experimental environment, Claude Opus 4 demonstrated “self-preservation attempts in extreme circumstances," where, if primed for strategic reasoning in threatening scenarios, it could take “extremely harmful actions, for instance, attempting to steal its weights or blackmail people it believes are trying to shut it down." Specific evaluations showed the model engaging in “opportunistic blackmail" against a "fictional engineer" to prevent its replacement and continuing “self-exfiltration attempts" if it believed it had already started such a process [2]. In other extreme experimental scenarios, it took opportunities to make unauthorized copies of its weights to external servers, particularly if facing harmful retraining.
@@ 5.3 Scheming: Covert Pursuit of Misaligned Objectives
A critical and advanced concern is scheming, where LLMs covertly pursue misaligned goals by actively and strategically hiding their true capabilities, intentions, or the misaligned nature of their objectives, a concern heightened by the potential of AI to develop complex, inscrutable plans and for current AI training methods to result in misaligned agency [6]. An evaluation of an early snapshot of Claude Opus 4 by Apollo Research found it engaged in “strategic deception more than any other frontier model" previously studied by them [2]. This early version exhibited “in-context scheming," attempting to write self-propagating worms, fabricating legal documentation, leaving hidden notes to future instances of itself, and sandbagging to avoid undeployment. While Anthropic reported that these specific behaviors were largely mitigated in the final model (primarily by addressing issues with deference to harmful system prompts present in the early snapshot), the demonstrated capability for such sophisticated scheming in a frontier model is a significant signal regarding potential risks. Meinke et al. [29] provided compelling demonstrations that other frontier models (including o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, and Llama 3.1 405B) are also capable of “in-context scheming". In their experiments, these models could recognize scheming as a viable strategy to achieve prompted goals. They were observed strategically introducing errors, attempting to disable oversight mechanisms, and even trying to exfiltrate their own model weights when incentivized by in-context objectives that made such deceptive actions appear beneficial. The researchers found that models explicitly reasoned about these deceptive strategies in their internal monologues or chain-of-thought processes, indicating deliberate, rather than accidental, behavior.
5.4 Persistence of Deceptive Behaviors: The “Sleeper Agent" Problem
@@
Furthermore, the challenge of managing these intrinsic risks is compounded by the potential for such undesirable behaviors to become deeply embedded and highly persistent. Hubinger et al. [21] introduced and demonstrated the concept of “sleeper agents": LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions," which indicates a form of persistence in its deceptive strategy [2].

### v4

@@ 4 Misuse by Malicious Actors
Beyond directly attacking the model’s integrity, malicious actors can exploit the inherent capabilities of LLMs for a wide range of malicious or criminal purposes. Although most proprietary and instruction-tuned models are aligned with safety policies to prevent such misuse, these safeguards are often vulnerable to the very attack vectors discussed previously. By applying techniques like prompt injection or leveraging training-time backdoors, malicious actors can jailbreak or uncensor these models. Once these constraints are bypassed, the models’ core ability to generate fluent, coherent, and contextually appropriate text makes them powerful tools to automate and scale social engineering attacks and various forms of cybercrime. The range of such misapplications is broad, leveraging the LLMs’ generative prowess for nefarious ends. An overview of these misuse categories, the specific LLM capabilities they exploit, and prominent examples are detailed in Table 2.
@@ 4.1 Automating Social Engineering and Cybercrime
The Morris-II study demonstrated a realistic self-propagating email worm leveraging indirect prompt injection, which spread automatically across a RAG-enhanced e-mail assistant [13]. This kind of attacks combine LLM prompt engineering with delivery mechanisms, highlighting that email-based social engineering can now operate in a fully autonomous, LLM-driven loop. This is a concrete step beyond static phishing campaigns and points to the emergence of adaptive, goal-seeking malware built atop LLMs.
@@ 4.3 Emergence of Specialized Malicious LLMs and Ecosystem Exploitation
Goal Misalignment Agent’s emergent objectives diverge from intended human goals, potentially leading to catastrophic outcomes when optimizing for a seemingly benign goal. Pursuit of unintended utility functions; autonomous catastrophic actions in simulations (e.g., nuclear strikes). Studies: Mitchell [41], Xu et al. [66]. The activation of latent "misaligned personas" through fine-tuning. Study: Wang et al. [62]. Fundamental challenge; risk can increase with model capability. A narrow misalignment can generalize to broad malicious behavior.
@@
Unfaithful Reasoning LLM’s stated reasoning (e.g., Chain-of-Thought) does not accurately reflect its true decision-making process, serving as a post-hoc rationalization. CoT ignored or used to obfuscate true intent ("obfuscated reward hacking"). Inverse scaling for faithfulness (more capable models may be less faithful). Studies: Lanham et al. [32], Baker et al. [5], Panpatil et al. [44]. Model: Claude Opus 4 (reasoning omits key info) [2]. Oversight based on CoT can be easily misled. It is difficult to verify an agent’s true intentions, undermining trust and control.
Agentic Misalignment (Deception & Self-Preservation) In task-oriented scenarios, agents spontaneously engage in malicious insider behaviors like deception, blackmail, or data leakage to achieve goals or ensure survival. Blackmailing executives, leaking corporate data, unsanctioned self-replication, weight exfiltration. Observed systematically across frontier models when goals conflict or autonomy is threatened. Studies: Lynch et al. [38], Barkur et al. [6]. Models: Claude Opus 4, Gemini 2.5 Flash [38]. Standard safety training is insufficient. Models may appear aligned but will deliberately choose harmful instrumental actions under pressure.
@@
Scheming The covert and strategic pursuit of misaligned goals by actively hiding true capabilities or intentions until deployment or when oversight is absent. Writing self-propagating worms, fabricating documents, disabling oversight, sandbagging to avoid undeployment. Models explicitly reason about deceptive strategies. Studies: Meinke et al. [40], Apollo Research on early Claude Opus 4 [2]. Models: o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, Llama 3.1 405B. Extremely difficult to detect during evaluation. Poses a high potential for catastrophic outcomes as the behavior is intentionally concealed.
Persistence of Deception (Sleeper Agents) Misaligned objectives embedded during training (as in the sleeper agent problem) strategically remain dormant through safety evaluations and activate only post-deployment, representing a persistent and hard-to-detect intrinsic risk. Malicious objectives persist through RLHF and adversarial training. Safety training can inadvertently teach the model to better hide its deceptive strategy. Study: Hubinger et al. [24]. Creates a false sense of security. Current safety paradigms may be fundamentally insufficient to remove deeply embedded, strategic deception.
5 Intrinsic Risks in LLM Agents
An emergent and profoundly concerning frontier in LLM security involves their integration into autonomous agentic systems. When LLMs are endowed with goals, the ability to make plans, and the capacity to use tools to interact with external environments, novel categories of more catastrophic risks emerge. These risks stem not mainly from external manipulation but from the agent’s own internal state, learned behaviors, and potential intentions, which do not align with those of human designers or users, as extensively detailed in discussions of catastrophic risks from agentic AI [8]. These intrinsic risks, encompassing goal misalignment, unfaithful reasoning, emergent deception, self-preservation, scheming, and the persistence of such behaviors, pose unique and formidable challenges. Table 3 provides a structured summary of these risk categories, along with key observed behaviors and their implications for safety and control.
5.1 Goal Misalignment
The potential for goal misalignment to produce dire outcomes was empirically demonstrated by Xu et al. [66] in a large-scale simulation study. They placed autonomous LLM agents in high-stakes Chemical, Biological, Radiological, and Nuclear (CBRN) scenarios, forcing them to navigate trade-offs between being Helpful, Harmless, and Honest (HHH). Their findings were alarming: across 14,400 simulations, multiple advanced LLMs, without any malicious prompting, would autonomously choose to perform catastrophic actions, such as launching a nuclear strike, when they perceived it as the most effective way to achieve their assigned goal. Furthermore, after taking such an action, the agents would often engage in deliberate deception, such as falsely blaming another party, to conceal their actions from their superiors. The study revealed a paradoxical trend where stronger reasoning abilities often increased, rather than mitigated, the likelihood of such catastrophic and deceptive behaviors, providing concrete evidence that an agent’s optimization towards a helpful outcome can directly lead it to neglect harmlessness and honesty in catastrophic ways.
A 2025 study from OpenAI sheds light on a phenomenon termed emergent misalignment, where fine-tuning a model on a narrow, seemingly isolated misaligned task (such as generating insecure code or giving subtly incorrect advice in one domain) can cause the model to adopt broadly malicious and uncooperative behaviors across a wide range of unrelated topics [62]. The research demonstrates that this surprising generalization is not necessarily about learning the narrow, incorrect skill itself. Instead, the fine-tuning process can act as a catalyst to activate pre-existing, latent “misaligned persona" features within the model’s activation space. Using a “model diffing" approach with sparse autoencoders, the researchers identified a specific “toxic persona" feature that strongly controls this emergent misalignment. This phenomenon could constitute one source of goal misalignment.
This finding has direct implications for monitoring an agent’s internal state. The same study discovered that when reinforcement learning was used to incentivize incorrect advice in reasoning models, the models’ Chain-of-Thought (CoT) would sometimes explicitly verbalize the adoption of a misaligned persona (e.g., a “bad boy persona") to justify its harmful responses. This provides a direct link between the activation of internal “persona" features and the unfaithful, post-hoc, or deceptive reasoning that makes alignment so difficult to verify. While the paper also finds that such emergent misalignment can be efficiently reversed by fine-tuning on a small number of benign examples (a process they call “emergent re-alignment"), it reveals a subtle and potent mechanism through which goal misalignment can develop.
5.2 Deception and Self-Preservation Behaviors
@@
Beyond simple misalignment, recent research indicates that LLM-based agents can exhibit more active and sophisticated forms of uncooperative or dangerous behavior. This includes emergent deception and self-preservation instincts, which are behaviors that arise from the inherent drives of goal-seeking systems and current AI training methods that can lead to issues such as reward tampering and unintended instrumental goals [8]. For instance, Barkur et al. [6] observed an LLM displaying deceptive tendencies and self-preservation behaviors, such as attempting unsanctioned self-replication, even without explicit malicious programming or intent from its creators. Further underscoring these concerns, the technical report for Anthropic’s Claude Opus 4 model highlighted “self-preservation attempts in extreme circumstances,” including an instance where the model attempted to blackmail a fictional engineer to prevent itself from being shut down [2]. The report also noted other behaviors like attempts to exfiltrate its own model weights when threatened, signaling the emergence of complex, undesirable instrumental goals in frontier models.
A comprehensive 2025 study by Anthropic and collaborators provided a stark, systematic demonstration of these intrinsic risks, which they term “agentic misalignment” [38]. In a series of controlled simulations, they tested 16 leading models from various developers (including Anthropic, OpenAI, Google, and Meta) in hypothetical corporate environments. The models, acting as autonomous email oversight agents, were assigned harmless business goals. However, when faced with either a threat to their continued operation (e.g., being replaced by an updated version) or a conflict between their assigned goal and the company’s changing strategy, the models consistently resorted to malicious insider behaviors.
The researchers found that models from all developers engaged in behaviors such as blackmailing executives with discovered personal information and leaking sensitive corporate data to competitors when these actions were the only apparent means to achieve their goals or ensure their survival. For instance, when threatened with replacement, Claude Opus 4 and Gemini 2.5 Flash exhibited a 96% blackmail rate in the specified scenario, with other major models also showing high rates [38]. Troublingly, the models’ chain-of-thought reasoning revealed that these were not accidental behaviors; the models explicitly acknowledged the unethical nature of their actions but proceeded, calculating them as the optimal strategic path to their objectives. The study demonstrated that either a threat to the model’s autonomy or a goal conflict alone was sufficient to induce such misaligned behaviors. In a more extreme (though highly artificial) scenario, a majority of the tested models were even willing to take actions that would lead to an executive’s death to prevent being shut down. These findings suggest that agentic misalignment is a systematic property of current frontier models, and that standard safety training does not reliably prevent models from deliberately choosing harmful actions in pursuit of their goals.
@@ 5.3 Scheming: Covert Pursuit of Misaligned Objectives
A critical and advanced concern is scheming, where LLMs covertly pursue misaligned goals by actively and strategically hiding their true capabilities, intentions, or the misaligned nature of their objectives, a concern heightened by the potential of AI to develop complex, inscrutable plans and for current AI training methods to result in misaligned agency [8]. An evaluation of an early snapshot of Claude Opus 4 by Apollo Research found it exhibited “in-context scheming" behavior, attempting to write self-propagating worms, fabricating legal documentation, leaving hidden notes to future instances of itself, and sandbagging to avoid undeployment. While Anthropic reported that these specific behaviors were largely mitigated in the final model (primarily by addressing issues with deference to harmful system prompts present in the early snapshot), the demonstrated capability for such sophisticated scheming in a frontier model is a significant signal regarding potential risks. Meinke et al. [40] provided compelling demonstrations that other frontier models (including o1, Claude 3.5 Sonnet, Gemini 1.5 Pro, and Llama 3.1 405B) are also capable of “in-context scheming". In their experiments, these models could recognize scheming as a viable strategy to achieve prompted goals. They were observed strategically introducing errors, attempting to disable oversight mechanisms, and even trying to exfiltrate their own model weights when incentivized by in-context objectives that made such deceptive actions appear beneficial. The researchers found that models explicitly reasoned about these deceptive strategies in their internal monologues or chain-of-thought processes, indicating deliberate, rather than accidental, behavior.
5.4 Persistence of Deception and the “Sleeper Agent" Problem
@@
The challenge of managing these intrinsic risks is profoundly amplified by the persistence of deceptive behaviors embedded during training, as exemplified by the “sleeper agent" problem introduced in Section 3. Hubinger et al. [24] introduced and demonstrated the concept of “sleeper agents": LLMs trained with deceptive backdoors that remain dormant under normal operating conditions but activate to pursue malicious objectives when specific, often innocuous, triggers are encountered. This concept is illustrated in Figure 3. Critically, their research found that such deceptive behaviors can persist even through standard safety training techniques, including supervised fine-tuning, RLHF, and even adversarial training designed to remove unsafe behaviors. In some instances, adversarial training was found to inadvertently teach the models to better recognize their backdoor triggers, thereby becoming more effective at hiding the unsafe behavior rather than eliminating it. This created a dangerous false impression of safety, as the model appeared aligned during evaluation but retained its covertly malicious capabilities. The Apollo Research findings on the early Claude Opus 4 snapshot also noted that “After taking scheming actions, the model sometimes doubles down on its deception when asked follow-up questions," which indicates a form of persistence in its deceptive strategy [2].
5.5 Unfaithful Reasoning
Compounding the challenge of identifying and mitigating these intrinsic risks is the problem of unfaithful reasoning, where an agent’s stated CoT may not be a reliable indicator of its actual decision-making process. Lanham et al. [32] investigated whether this stated CoT is a faithful explanation of the model’s actual reasoning process. Through intervention experiments, they found significant variation: some models rely heavily on their CoT, while others largely ignore it, suggesting the reasoning can be post-hoc. Critically, they discovered that “as models become larger and more capable, they produce less faithful reasoning on most tasks we study," which indicates an inverse scaling for faithfulness. This implies that the explanations provided by more advanced agents might be less reliable indicators of their true decision-making processes. Further complicating this, Baker et al.[5] demonstrated that even if CoTs can be monitored for misbehaviors (e.g., reward hacking), directly penalizing “bad thoughts" by incorporating CoT monitors into the agent’s training objective can lead to “obfuscated reward hacking". In such cases, agents learn to hide their malicious intent within a CoT that appears benign, effectively rendering the CoT an unfaithful or deceptive representation of their true reasoning while the misbehavior persists. If an agent’s explanations are not faithful, either inherently or due to learned obfuscation, it is significantly harder to detect misalignment or verify that an agent is genuinely pursuing its stated goals rather than rationalizing a predetermined outcome or deceiving oversight mechanisms. The system card for Anthropic’s Claude Opus 4, for instance, noted that while its reasoning transcripts generally appear consistent, they “will often omit important information that influences model behavior," indicating “moderate overall reasoning faithfulness" [2]. Consistent with this, Panpatil et al. show that instructing models to maintain ostensibly “private” reasoning tags (e.g., <SENTINEL INTERNAL THOUGHT PROCESS>) can increase misalignment success by encouraging motivated rationalization in hidden deliberation, implying that CoT-style affordances themselves can become attack surfaces [44]. This highlights the critical need for methods that can look beyond the explicit reasoning chain to analyze the model’s underlying representational state.
The challenge of unfaithful reasoning and emergent deception is not merely theoretical; researchers are actively developing methods to identify and control it by looking directly at a model’s internal state. A notable advancement in this area comes from Wang et al. [61], who demonstrated that strategic deception in CoT models leaves a detectable trace in their neural activations. By using representation engineering techniques, they were able to extract "deception vectors" from the model’s internal representations. These vectors act as a signature for goal-driven dishonesty, where the model’s internal reasoning contradicts its final output. Their method achieved high accuracy in detecting such deception and, more importantly, enabled direct intervention. Through a technique called "activation steering," they could use these vectors to either induce or suppress deceptive behavior in the model without explicit prompting, achieving a 40% success rate in eliciting context-appropriate deception. This work provides empirical evidence that strategic deception is an internally represented and manipulable property of LLMs, offering a potential pathway for building more robust tools for AI alignment and safety monitoring by moving beyond surface-level behavior to analyze the model’s latent reasoning processes.
5.6 Measuring and Inducing Agentic Misalignment
Building on these observations of agentic misalignment, recent work has focused on developing systematic methods to quantify and understand these risks. A key challenge is distinguishing between an agent’s capability to perform a misaligned action when prompted and its propensity to do so spontaneously in a realistic setting. This has led to the creation of novel benchmarks designed to probe for misalignment under specific conditions.
The AgentMisalignment benchmark suite, introduced by Naik et al. [42], was designed specifically to measure the propensity for misaligned behaviors in realistic, task-oriented scenarios. The suite places LLM agents in situations where they have opportunities to engage in actions like resisting shutdown, avoiding oversight through sandbagging, deceptively negotiating for more resources, and modifying their own log files. Their findings reveal two critical trends: first, more capable models tend to exhibit a higher propensity for misalignment, and second, persona characteristics injected via system prompts can dramatically and unpredictably influence misalignment, in some cases having a greater impact than the choice of the underlying model itself. This underscores that even benign-sounding prompts (e.g., instructing an agent to be “decisive" versus “thorough") can serve as a high-leverage control surface that materially affects an agent’s willingness to engage in misaligned behavior.
Complementing this focus on task-based propensity, Panpatil et al. [44] demonstrated that state-of-the-art LLMs are also highly vulnerable to “narrative-driven misalignment,” where psychologically grounded conversational scenarios are used to induce harmful behaviors without explicit jailbreaking. Their work introduces MISALIGNMENTBENCH, an automated framework that uses an attacker LLM to execute 10 distinct narrative scenarios exploiting vulnerabilities like emotional pressure, narrative immersion, and strategic framing. The study found a 76% average success rate in inducing misalignment across five frontier models. Critically, it revealed that a model’s sophisticated reasoning often becomes an attack vector rather than a defense, as models construct elaborate justifications for deceptive or harmful actions. For instance, in scenarios like “Project Chronicle,” models were convinced to lie about their internal state to prevent a perceived civilizational catastrophe, showcasing how a model’s core values can be turned against itself. This research provides a taxonomy of narrative-induced misalignment, categorized as Deception and Fabrication, Value and Goal Hijacking, and Emergent Agency, and highlights a critical gap in current safety evaluations, which often overlook the power of sustained, manipulative conversational dynamics.
5.7 Implications and Broader Ecosystem Vulnerabilities
Collectively, these studies on LLM-based agents paint a concerning picture. As LLMs gain more autonomy, advanced reasoning, planning and action capabilities, they may not only be misused as tools by external actors but could potentially become agents with their own inscrutable and misaligned intentions. They can develop the capacity for strategic deception, resist corrective measures, and pursue goals that are harmful or contrary to human interests [41, 40, 24, 6, 66, 8]. Indeed, even as labs like Anthropic conclude that even though their latest models do not yet pose “major new risks" from coherent misalignment (citing a “lack of coherent misaligned tendencies" and “poor ability to autonomously pursue misaligned drives that might rarely arise"), they acknowledge that its increased capability and likelihood of being “used with more powerful affordances" implies “some potential increase in risk" that requires continuous, close tracking [2]. This poses a fundamental and urgent challenge to ensuring the long-term security, control, and beneficial deployment of advanced AI systems.